# Shortcuts to adiabaticity for coherent atom transport in an adjustable family of two-dimensional optical lattices

### v1

@@ VII Summary and Conclusions
To conclude, in this paper we investigated coherent atom transport in an adjustable family of two-dimensional optical lattices. We first specified concrete arrangements of acousto-optic modulators that give rise to the moving-lattice effect in these systems in an arbitrary direction. We subsequently designed the corresponding dynamical-lattice trajectories using shortcuts to adiabaticity (STA). Through numerical propagation of the relevant time-dependent Schrödinger equation – using the previously computed STA-based dynamical-lattice trajectories – we calculated the efficiency ofhe resulting coherent atom transport as described by the atom-transport fidelity.
We demonstrated that the STA approach enables fast atom transport in the considered family of optical lattices, for a wide range of lattice depths and transport directions. In addition, we analyzed the implications of the atom-transport processes for neutral-atom quantum computing based on collisional two-qubit entangling gates.
By contrast to previous theoretical investigations of coherent atom transport, based mostly on simplified scenarios, in the present work we addressed this phenomenon based on the full, anharmonic optical-lattice potential. Thus, the obtained results can be corroborated in future transport experiments in such lattices. Furthermore, our study also complements well the recent works on quantum-state control of Bose-Einstein condensates in optical lattices Amr . Finally, this work may provide the impetus for further experimental implementations of STA-based control schemes for coherent atom transport in optically-trapped neutral-atom systems, the first one of which has been undertaken quite recently Hwa .

### v2

@@ VII Summary and Conclusions
To conclude, in this paper we investigated coherent atom transport in an adjustable family of two-dimensional optical lattices which includes square, honeycomb, dimerized, and 1D-chains lattices as its special cases. We first specified concrete arrangements of acousto-optic modulators that give rise to the dynamical(moving)-lattice effect in these systems in an arbitrary direction. We subsequently designed the corresponding dynamical-lattice trajectories using both shortcuts to adiabaticity (STA) in the form of inverse engineering based on a Lewis-Riesenfeld type dynamical invariant, as well as their enhanced version (eSTA). Through the numerical propagation of the relevant time-dependent Schrödinger equation – using the previously obtained STA and eSTA dynamical-lattice trajectories – we quantified the efficiency of the resulting coherent atom transport as described by the atom-transport fidelity.
By computing transport fidelities for wide ranges of values of relevant parameters (lattice depths, transport distances, etc.) and different transport directions for four different lattice configurations, we showed that – with the notable exception of the dimerized lattice – the eSTA method consistently outperforms its STA counterpart; this superiority of the eSTA approach is manifested both in terms of the achievable transport times and the robustness of the resulting transport against small variations of optical-lattice depths. For the sake of completeness, we also demonstrated that our adopted eSTA approach to atom transport is far superior to alternative approaches to modeling atom transport in optical lattices, more precisely the approaches based on ad-hoc velocity profiles.
By contrast to previous theoretical investigations of coherent atom transport, based mostly on simplified scenarios, in the present work we addressed this phenomenon based on the full, anharmonic optical-lattice potential. Thus, the obtained results can be corroborated in future transport experiments in such lattices. Furthermore, our study also complements well the recent works on quantum-state control of Bose-Einstein condensates in optical lattices Amr . Finally, this work may provide the impetus for further experimental implementations of STA-based control schemes for coherent atom transport in optically-trapped neutral-atom systems (including those based on the eSTA approach), the first one of which has been undertaken quite recently Hwa .